OpenStudy (anonymous):

Find the general antiderivative of f(t)=6t

geerky42 (geerky42):

\[\int cx^n \mathrm d x =c\int x^n\mathrm dx= \dfrac{cx^{n+1}}{n+1}\]

geerky42 (geerky42):

I forgot constant lol.

OpenStudy (anonymous):

ok lol im still lost

geerky42 (geerky42):

You are familiar with some antiderivative rules? @mike95

OpenStudy (anonymous):

not really

geerky42 (geerky42):

What have you learned so far?

OpenStudy (anonymous):

im just not familiar with antiderivatives at all

OpenStudy (anonymous):

so if you could walk me through this problem i would appreciate it

geerky42 (geerky42):

well, I "learned" antiderivative by memorized rules lol... Can you help @jim_thompson5910 ?

OpenStudy (anonymous):

do you know where i could find a chart of these rules then

geerky42 (geerky42):

You could google for it. Here is one of them: http://www.wyzant.com/resources/lessons/math/calculus/integration/antiderivatives

geerky42 (geerky42):

better one for this problem: https://math.la.asu.edu/~andrzej/teachold/210Sum04/teach/rulesant.pdf

OpenStudy (anonymous):

so tell me if i did this right...

OpenStudy (anonymous):

im gonna change it to f(x) because its easier for me

OpenStudy (anonymous):

f(x)=6x is our given

OpenStudy (anonymous):

we use the power rule and get 6x^2/2+c

OpenStudy (anonymous):

which results in 3x^2+c

geerky42 (geerky42):

That's right.

OpenStudy (anonymous):

thanks

geerky42 (geerky42):

no problem

OpenStudy (anonymous):

another one?

OpenStudy (anonymous):

@geerky42 f(x)=x^2-4x+7

OpenStudy (anonymous):

x^3/3-2x^2+7x+c

geerky42 (geerky42):

Correct.

OpenStudy (anonymous):

ok im getting the hang ouf it thanks again pal.